### v1

# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

Abstract
Online polarization poses a growing challenge for democratic discourse, yet most computational social science research remains monolingual, culturally narrow, or event-specific. We introduce POLAR, a multilingual, multicultural, and multievent dataset with over 23k instances in seven languages from diverse online platforms and real-world events. Polarization is annotated along three axes: presence, type, and manifestation, using a variety of annotation platforms adapted to each cultural context. We conduct two main experiments: (1) we fine-tune six multilingual pretrained language models in both monolingual and cross-lingual setups; and (2) we evaluate a range of open and closed large language models (LLMs) in few-shot and zero-shot scenarios. Results show that while most models perform well on binary polarization detection, they achieve substantially lower scores when predicting polarization types and manifestations. These findings highlight the complex, highly contextual nature of polarization and the need for robust, adaptable approaches in NLP and computational social science. All resources will be released to support further research and effective mitigation of digital polarization globally.
POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization
Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University,4Imperial College London 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
Figure 1: Pipeline for POLAR construction: multi-platform data curation in 7 languages, annotation workflow with quality control, and benchmarking.
1 Introduction
Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (Waller and Anderson, 2021; Iandoli et al., 2021). It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (Garimella, 2018). Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence. As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
Despite growing attention, computational approaches to polarization suffer from major limitations. First, most existing datasets focus on English or high-resource languages, reflecting a widespread trend across NLP tasks that ignores the rich diversity of linguistic and sociocultural contexts in which polarization manifests. Second, current benchmarks are often event-specific or monodomain, such as U.S. elections or Western political debates, limiting their generalizability. Third, the conceptualization of polarization in NLP has largely been binary or topic-focused, overlooking the multifaceted ways in which polarization is expressed through vilification, dehumanization, stereotyping, or other rhetorical tactics.
To address these gaps, we introduce POLAR a novel multilingual, multicultural, and multievent dataset for fine-grained polarization detection. It spans seven languages across diverse regions, including low-resource languages such as Amharic and Hausa. Our data is sourced from various platforms (e.g., Twitter/X, Facebook, BlueSky, Reddit, and local news outlets), reflecting authentic, event-driven discourse ranging from armed conflict (e.g., the Tigray War) to social justice movements (e.g., abortion rights, migration crises).
Unlike prior work, POLAR supports three complementary tasks:
  1. 1.
    Binary Polarization Detection: Is a message polarized or not?
  2. 2.
    Polarization Type Classification: What social dimension is targeted (e.g., political, religious, racial)?
  3. 3.
    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping, deindividuation, extreme language)?
We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context. Extensive qualitative analysis reveals how event salience, linguistic norms, and platform affordances shape polarization dynamics across languages. The complete pipeline can be found in Figure 1. We benchmark a range of multilingual language models (MLMs) and large language models (LLMs) under zero-shot, few-shot, and cross-lingual scenarios. Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages. Our contributions are as follows:
  • We release POLAR, the first large-scale, multilingual, fine-grained dataset for polarization detection across 7 languages and diverse global events.
  • We define a taxonomy of polarization types and manifestations, operationalized through a robust cross-lingual annotation protocol.
  • We provide comprehensive benchmarks using state-of-the-art MLMs and LLMs across multiple evaluation settings (monolingual, cross-lingual, and few-shot).
Figure 2: Languages represented in the dataset, covering diverse linguistic and regional contexts.
2 Related Work
Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (Waller and Anderson, 2021; Iandoli et al., 2021; Garimella, 2018). As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly. Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance. Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (Arora et al., 2022), and has highlighted its relationship with hate speech, fragmentation, and incivility (Bozdag and van den Hoven, 2020; Mathew et al., 2021).
A growing body of research has documented the role of online spaces in intensifying polarization across different regions (Kubin and von Sikorski, 2021; Barberá, 2020; Gitlin, 2016; Soares and Recuero, 2021). However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (Kubin and von Sikorski, 2021). This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities. Recent shared tasks on hate speech and toxicity (Basile and others, 2019; Mohammad and others, 2021; Pamungkas et al., 2020; Ousidhoum et al., 2024) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations. Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
3 POLAR Dataset Construction
3.1 Data Collection
We collected data from various online platforms, including X (formerly Twitter), Facebook, Reddit, Bluesky, Threads, and news/commentary forums. We use a dynamic keyword-driven strategy tailored for each language. Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events. Table 1 shows the languages covered, data splits and total number of instances annotated for each language.
The Amharic dataset focuses on the Northern Ethiopia or Tigray War in Ethiopia; Arabic texts cover a broad array of social and economic topics; English data centers on US elections and international conflicts; the German corpus features election discourse and migration debates; the Hausa subset captures religious and ideological discussions in West and Central Africa; Spanish content spans abortion, migration, and indigenous/gender rights; and Urdu texts reflect political and sectarian divides.
Language Source(s) Train Dev Test Total
Amharic Facebook, X 3500 500 1000 5000
Arabic Facebook, X, Threads, News 1482 212 424 2118
English X, BlueSky, News 2117 303 605 3025
German X, BlueSky, Reddit 2426 347 694 3467
Hausa Facebook , X 3893 557 1113 5563
Spanish X, BlueSky 1400 201 401 2002
Urdu X 1960 280 560 2800
Table 1: Dataset sources and split sizes for all three tasks as per each language.
3.2 Annotation Process
Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators.
Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language. Annotators were instructed to:
  • Identify whether a text is polarized;
  • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other);
  • If the text is classified as polarized, tag its manifestations (stereotyping, vilification, dehumanization, deindividuation, extreme language, lack of empathy, invalidation).
Multiple labels were allowed due to the conceptual and contextual overlap often observed in polarized content.
Annotator Recruitment and Annotation Process
We used both expert and crowd-sourced annotation strategies. Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific. To evaluate inter-annotator agreement, we report both Cohen’s Kappa and Fleiss’ Kappa. Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
Language Total Polarization Polarization Types Polarization Manifestations
gender/sexual political religious racial/ethnic other vilification extreme_language stereotype invalidation lack_of_empathy dehumanization
Amharic 5000 3753 29 3342 99 1297 189 2398 1527 2729 799 880 657
Arabic 2118 338 102 178 86 171 220 413 341 283 240 175 102
English 3025 1021 18 892 44 98 5 342 179 138 85 45 39
German 3467 1596 241 1981 448 782 653 1371 1001 1420 1258 1124 568
Hausa 5563 625 46 285 149 183 22 74 179 253 13 53 202
Spanish 2002 1057 254 742 571 521 405 798 622 843 205 599 307
Urdu 2800 1860 236 1533 557 423 126 1548 1409 979 670 632 501
Table 2: Number of samples labeled positive for each annotation task across languages. Labels are grouped by task: Polarization (Task 1), Polarization Types (Task 2), and Polarization Manifestations (Task 3).
  • Amharic: We used train annotators to annotated 5,000 samples in five batches using the POTATO annotation tool. Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control. The annotation quality achieved a Fleiss’ Kappa of 0.49, with full agreement in 69.62% of cases and a peak pairwise Cohen’s Kappa of 0.65.
  • Arabic: Three topic-aware annotators labeled 2,118 samples using POTATO. Final polarization distribution was 16% positive, with Cohen’s Kappa at 0.296. Annotations were revised to ensure thematic consistency.
  • German: Annotation was conducted through Prolific. Annotators were screened via surveys and underwent pilot testing. Final annotation used a locally hosted POTATO interface adapted for multilabel inputs. Inter-annotator agreement was moderate, with notable consistency on polarized cases.
  • Hausa: We used trained annotators and Label Studio with iterative feedback and internal consistency checks. Data focused on religion and ideology on social media.
  • Spanish and English: Amazon Mechanical Turk (MTurk) was used with strict quality filters, such as task completion history and approval ratings. Annotators received feedback and were contacted throughout for quality assurance. English annotation reached a Kappa of 0.52 (up from 0.31 after worker filtering); Spanish showed lower agreement with an average Kappa of 0.24.
  • Urdu: Two approaches were used: (1) Manual Annotation: 1,792 samples labeled in batches with average Cohen’s Kappa of 0.55 (batch 1: 0.79; batch 2: 0.56; batch 3: 0.39). Disagreements were resolved by a fifth annotator. (2) Prolific: 1,460 samples were annotated in three batches. Initial agreement was low (Fleiss’ Kappa = 0.05). Additional filters (education, region) raised agreement in later batches to 0.28 and 0.30. The final dataset contains 2,800 high-quality Urdu samples.
Figure 3: Percentage of polarization per language
(a) Types
(b) Manifestation
Figure 4: Annotation statistics for polarization classification, types, and manifestations
3.3 Dataset Characteristics
Polarization Prevalence: As shown in Figure 3, Amharic (75%) and Urdu (66%) has the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate. Spanish (53%), German (46%), and English (34%) fall in the mid-range. Arabic (12%) and Hausa (11%) reflect broader topical focus and possibly platform moderation effects. Hausa shows the lowest polarization rate despite covering sensitive religious and ideological topics, possibly due to moderation or cultural norms in expression. Datasets with acute sociopolitical events (e.g., Amharic, Urdu) show higher polarization. Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
Types of Polarization: Figure 4 show types of polarization. Political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%). Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German. Gender/sexual identity polarization is generally low but visible in Spanish (13%) and Urdu (8%). Thematic emphasis (e.g., migration, elections) shapes type distribution. Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
Polarization Manifestations: Figure 4 show types of Manifestation. Stereotyping and vilification are most frequent in Urdu, Amharic, German, and Spanish. Urdu’s vilification rate is highest (55%). Extreme language is strong in Urdu (50%), Spanish (31%), and Amharic (31%). Invalidation and dehumanization peak in German and Urdu. Conflict-heavy datasets (Urdu, Amharic) correlate with higher rates of hostile manifestations. Platform moderation and topic scope likely suppress such features in Arabic and Hausa.
Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization. The POLAR dataset (see Table 2 for detailed statistics of POLAR dataset) offers a robust benchmark for examining these effects at scale.
4 Experimentation and Results
4.1 Experimental Setup
To evaluate POLAR , we conducted baseline experiments on three polarization detection tasks:
  • Task 1: Binary classification - determining whether a text is polarized.
  • Task 2: Multi-label classification - identifying polarization types (e.g., political, religious, ethnic).
  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., incivility, stereotyping, dehumanization).
For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table 1. We pursued two main experimental paradigms:
  1. 1.
    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023) for monolingual and crosslingual experiments.
  2. 2.
    Evaluating Large Language Models (LLMs) in Zero- and Few-shot Settings: We tested models including QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, and GPT4o/mini.
4.2 Results and Analysis
Monolingual Cosslingual
Task Lang. InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa
1. Polarization Amharic 77.96 78.92 81.93 80.64 53.71 74.98 23.26 49.44 5.63 0.00 0.00 83.21
Arabic 58.39 61.94 67.19 70.42 52.48 54.68 29.68 35.87 27.76 2.78 37.79 25.63
English 72.09 69.84 75.43 76.08 72.77 74.94 1.90 51.68 53.41 46.89 33.44 53.45
German 29.75 63.46 67.50 67.78 58.45 64.38 0.62 51.97 64.59 41.87 46.05 25.94
Hausa 59.57 66.41 67.43 66.41 59.32 65.62 8.21 22.92 6.47 1.50 6.25 25.13
Spanish 38.69 64.04 70.98 57.00 59.31 54.16 45.70 68.66 67.26 21.09 68.34 62.17
Urdu 1.07 66.35 79.95 43.91 65.63 60.68 1.60 30.36 0.00 0.00 3.68 68.67
2. Types Amharic 24.22 38.13 43.65 25.71 17.93 25.00 11.56 20.65 2.49 0.00 0.00 26.76
Arabic 22.97 40.23 42.12 37.58 27.53 35.22 11.49 23.73 8.78 2.11 12.47 7.93
English 16.04 23.25 31.38 24.10 21.07 17.70 10.84 19.69 16.42 12.22 11.70 3.93
German 13.52 58.58 61.19 58.56 54.13 40.64 12.03 35.86 29.59 9.72 23.98 11.88
Hausa 18.56 19.14 17.38 18.09 19.61 18.87 3.20 9.97 3.91 1.39 5.78 5.90
Spanish 43.26 66.07 67.76 58.07 57.94 43.40 7.89 47.06 15.38 0.43 32.02 14.60
Urdu 27.14 51.94 51.60 45.38 38.02 33.13 3.77 13.62 6.33 0.00 3.94 20.30
3. Manifestations Amharic 43.52 47.56 47.63 43.17 33.18 43.29 15.57 27.07 8.64 0.00 0.00 43.58
Arabic 40.05 51.55 52.52 55.61 42.18 47.73 16.56 30.68 22.14 0.00 19.57 17.24
English 14.40 15.01 19.39 18.61 18.60 15.15 7.16 10.16 10.05 5.62 10.69 8.85
German 38.49 49.88 52.74 51.70 46.85 51.91 2.38 36.12 27.05 0.00 23.38 12.93
Hausa 19.23 20.04 19.18 18.89 18.74 18.93 5.74 5.86 3.77 3.12 6.19 6.43
Spanish 38.94 50.00 51.04 45.02 45.17 35.09 2.34 40.63 11.83 0.40 35.99 23.56
Urdu 34.26 52.20 53.64 41.32 45.90 47.01 2.10 19.16 11.54 0.00 1.88 48.58
Table 3: Average F1-Macro for all three tasks. In the monolingual settings, we train and evaluate the model for each language separately. In the crosslingual setting, we train on all languages within a language family (AfroAsian: Amharic, Arabic, Hausa, and Urdu), Germanic: English, German, and Spanish) except the target language, and evaluate on the test set of the target language. The best performance scores are highlighted in blue and orange, respectively.
4.2.1 Monolingual Setup Results
To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023). Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
For Task 1, binary polarization classification, RemBERT consistently achieved the highest macro F1-scores for most languages, while XLM-R performed best in Arabic, English, and German. In the multi-label polarization types classification (Task 2), RemBERT again provided the top scores for Spanish, except for Hausa and Urdu, where mBERT and LaBSE performed best, respectively. For Task 3, multi-label polarization manifestations classification, the results were generally lower across all models, with RemBERT delivering the highest scores for most languages, except for Arabic and Hausa.
4.2.2 Cross-Lingual MLM Results
For cross-lingual transfer experiments, each MLM was trained on all other languages within the same language family (Afro-Asian or Germanic), excluding the target language, and evaluated on the target language’s test set for the same three tasks.
In the cross-lingual setup for polarization classification (Task 1), mDeBERTa achieved the highest F1-macro for Amharic, English, Hausa, and Urdu. For Arabic, German, and Spanish, mBERT, RemBERT, and LaBSE performed slightly better than the other models, respectively.
For Task 2, polarization types classification, performance was generally lower as expected in the cross-lingual scenario. LaBSE yielded the best results for four languages, while mDeBERTa performed better for Amharic and Urdu.
Similarly, for Task 3, polarization manifestation classification, the results were quite low overall, but LaBSE and mDeBERTa showed relatively better performance compared to the other models.
This suggests that general polarization detection transfers more easily across languages than the more nuanced identification of polarization types and manifestations, which are culturally and linguistically specific. Effective cross-lingual transfer likely requires aligned annotation schemes and fine-tuning strategies sensitive to these differences.
4.2.3 Zero and Few-shot LLM Performance
We evaluated five state-of-the-art LLMs (QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, GPT4o/mini) in a zero- and few-shot setting across languages and for the first ask. Larger models such as GPT4o and GPT4o-mini outperformed smaller ones, especially in capturing subtle polarization cues.
GPT 4o GPT 4o-mini Mistral-7B Llama-3.1-8B Qwen3-8B
Zero-shot Few-shot Zero-shot Few-shot Zero-shot Zero-shot Zero-shot
Amharic 71.30 68.74 55.40 62.59 71.92 36.32 58.09
Arabic 71.51 79.65 58.21 75.95 33.65 34.82 54.54
English 75.46 79.56 64.78 80.93 50.34 56.75 72.07
German 72.04 71.12 63.03 72.44 52.63 56.22 61.67
Hausa 34.73 44.87 28.21 44.26 23.42 18.51 25.06
Spanish 72.69 65.76 62.51 69.00 56.31 41.75 61.11
Urdu 72.25 78.83 59.19 74.66 70.47 67.26 69.63
Table 4: F1-Macro resulting from the zero- and few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
5 Discussion
Our multilingual and cross-cultural investigation into online polarization reveals that polarization is a deeply contextual and event-driven phenomenon. Its prevalence and intensity are closely tied to the sociopolitical climate, public discourse, and digital environment of each region. In societies experiencing major crises, such as the war in Ethiopia, religious tensions in Pakistan, and international conflicts like the wars in Ukraine and Gaza that influence the US election, we find polarization to be especially amplified on online platforms. Conversely, broader topical selection, moderation, or more diffuse societal tensions are associated with noticeably lower polarization.
The types of polarization evident in online discourse are shaped by the most salient local issues: political (Ethiopia, US, German), religious (Pakistan), ethnic (Ethiopia), or gender-based (Spain) divides tend to dominate where these subjects are at the forefront of public debate or conflict. This underscores the importance of accounting for each region’s unique sociopolitical history and cultural landscape when analyzing or intervening in digital polarization.
Polarization also manifests in varied rhetorical forms, including stereotyping, vilification, and exclusionary or hostile language. These patterns are particularly pronounced amid contentious issues and persistently divided communities, such as civil war, election, migration issues, religious divide, and so on.
From a computational perspective, our results demonstrate both the feasibility and the challenges of detecting polarization in multilingual settings. Binary polarization classification is relatively tractable for well-resourced languages and clear-cut contexts, especially when leveraging state-of-the-art multilingual models (Sec. 4. However, classifying the specific types and manifestations of polarization introduces significantly greater ambiguity and requires a much deeper, context-sensitive understanding. Performance on these more nuanced tasks remains limited. This highlights important directions for future work, including the integration of cultural signals and context or event understanding, the development of more robust multilingual embeddings, and the pursuit of consistent annotation (with a great deal of annotator training and follow up, specially for the crowdsourcing annotations). The wide variability across languages, contexts, and events further affirms the necessity of designing language, culture, and event specific approaches, particularly for underrepresented regions in global NLP research.
Taken together, these findings highlight that effective detection and mitigation of online polarization must move beyond generic or monocultural solutions. There is a pressing need for context-aware, adaptable methodologies that acknowledge both the universal and the local characteristics of polarized discourse, ensuring relevance and effectiveness across the world’s diverse digital landscapes.
6 Conclusion
In this study, we introduced POLAR, the first multilingual, multicultural, and multievent dataset for benchmarking online polarization. Our findings showed that polarization was deeply context-dependent, manifesting in diverse types and rhetorical forms shaped by local sociopolitical dynamics. Theoretically, we provide new evidence that polarization is a multifaceted, cross-lingual phenomenon strongly influenced by language, culture, and current events, rather than being uniform across settings. On a practical level, our benchmark revealed that while binary detection of polarization was feasible with current models, accurately identifying specific types and manifestations remained a persistent challenge, especially for multilingual and low-resource contexts.
This work established a foundation for further computational studies of digital polarization and for developing more culturally robust moderation and intervention tools. Looking ahead, future research should integrate richer contextual and cultural signals into model architectures, refine annotation guidelines for cross-regional consistency, and expand evaluations to include additional languages, social domains, and further events. By releasing our data and benchmarks, we aimed to catalyze further innovation toward nuanced detection and effective mitigation of online polarization worldwide.
Limitations
While POLAR represents an important step toward multilingual, multicultural, and multievent polarization analysis, several limitations remain. First, annotator understanding - particularly in crowdsourced setups - was sometimes limited, potentially impacting label quality. We mitigated this through strict quality assurance methods, including control questions, pre-study surveys, and ongoing annotator assessment, but some variability in interpretation may persist.
Second, in-house annotation, while yielding higher consistency, sometimes introduced psychological challenges for annotators given the sensitive or hostile nature of polarized content. To address this, we provided detailed instructions and support resources to reduce stress and clarify expectations, but some emotional burden may have remained.
Third, our choice of models is not exhaustive. Although we included several leading multilingual models and both open and closed LLMs. Adding more language-specific models in the future could improve results, especially for monolingual scenarios.
Finally, for some of the languages in our benchmark, the available data size is still limited, which may constrain the generalizability of model training and evaluation for those cases. Future work should expand dataset size and diversity, and explore language- or region-specific model development to better support underrepresented contexts.
Ethics Statement
This research uses only publicly available, anonymized data and addresses sensitive topics around polarization in diverse cultures. All annotation was conducted by native speakers using culturally appropriate guidelines; annotators were informed of the project’s social good aims, possible distress, and could opt out anytime. Annotators received prompt and fair compensation above local wage standards or per Prolific’s requirements. Despite rigorous protocols, labeling polarization remains subjective; we encourage responsible, ethically grounded use of this resource and discourage misuse.