### v2

# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

Abstract
Online polarization poses a growing challenge for democratic discourse, yet most computational social science research remains monolingual, culturally narrow, or event-specific. We introduce POLAR, a multilingual, multicultural, and multievent dataset with over 23k instances in seven languages from diverse online platforms and real-world events. Polarization is annotated along three axes: presence, type, and manifestation, using a variety of annotation platforms adapted to each cultural context. We conduct two main experiments: (1) we fine-tune six multilingual pretrained language models in both monolingual and cross-lingual setups; and (2) we evaluate a range of open and closed large language models (LLMs) in few-shot and zero-shot scenarios. Results show that while most models perform well on binary polarization detection, they achieve substantially lower scores when predicting polarization types and manifestations. These findings highlight the complex, highly contextual nature of polarization and the need for robust, adaptable approaches in NLP and computational social science. All resources will be released to support further research and effective mitigation of digital polarization globally.
POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization
Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University, 4Imperial College London, 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
1 Introduction
Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (waller2021quantifying; iandoli2021impact). It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (garimella2018polarization). Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence. As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
Figure 1: Pipeline for POLAR construction: multi-platform data curation in 7 languages, annotation workflow with quality control, and benchmarking.
Despite growing attention, computational approaches to polarization suffer from major limitations. First, most existing datasets focus on English or high-resource languages, reflecting a widespread trend across NLP tasks that ignores the rich diversity of linguistic and sociocultural contexts in which polarization manifests. Second, current benchmarks are often event-specific or monodomain, such as U.S. elections or Western political debates, limiting their generalizability. Third, the conceptualization of polarization in NLP has largely been binary or topic-focused, overlooking the multifaceted ways in which polarization is expressed through vilification, dehumanization, stereotyping, or other rhetorical tactics.
To address these gaps, we introduce POLAR a novel multilingual, multicultural, and multievent dataset for fine-grained polarization detection. It spans seven languages across diverse regions, including low-resource languages such as Amharic and Hausa. Our data is sourced from various platforms (e.g., Twitter/X, Facebook, BlueSky, Reddit, and local news outlets), reflecting authentic, event-driven discourse ranging from armed conflict (e.g., the Tigray War) to social justice movements (e.g., abortion rights, migration crises).
Unlike prior work, POLAR supports three complementary tasks:
  1. 1.
    Binary Polarization Detection: Is a message polarized or not?
  2. 2.
    Polarization Type Classification: What social dimension is targeted (e.g., political, religious, racial)?
  3. 3.
    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping/deindividuation, vilification, dehumanization, extreme language)?
We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context. The complete pipeline can be found in Figure 1. We benchmark a range of multilingual language models (MLMs) and large language models (LLMs) under zero-shot, few-shot, and cross-lingual scenarios. Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages. Our contributions are as follows:
  • We release POLAR, the first large-scale, multilingual, fine-grained dataset for polarization detection across 7 languages and diverse global events.
  • We define a taxonomy of polarization types and manifestations, operationalized through a robust cross-lingual annotation protocol.
  • We provide comprehensive benchmarks using state-of-the-art MLMs and LLMs across multiple evaluation settings (monolingual, cross-lingual, and few-shot).
Figure 2: Languages represented in the dataset, covering diverse linguistic and regional contexts.
2 Related Work
Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (waller2021quantifying; iandoli2021impact; garimella2018polarization). As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly. Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance. Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (arora2022polarization), and has highlighted its relationship with hate speech, fragmentation, and incivility (Mathew2021HateXplain).
A growing body of research has documented the role of online spaces in intensifying polarization across different regions (kubin2021role; barbera2020social; gitlin2016outrage; soares2021hashtag). However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (kubin2021role). This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities. Recent shared tasks on hate speech and toxicity (basile2019semeval; pamungkas2020misogynistic; ousidhoum2024semeval) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations. Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
3 POLAR Dataset Construction
We aim to construct a multilingual, multi-event, and multi-platform social media dataset spanning multiple continents. The dataset is designed to cover Indo-European and Afro-Asiatic language families, with a particular focus on balancing linguistic diversity and resource availability. It includes high-resource languages such as English, Spanish, and German; medium-resource languages like Arabic and Urdu; and low-resource languages such as Hausa and Amharic. While several Indo-European languages are represented (e.g., English, Spanish, German, Urdu), the dataset also incorporates key Afro-Asiatic languages (e.g., Hausa, Arabic, Amharic) to ensure broader geographical and linguistic representation. The following section will discuss the dataset construction in detail.
3.1 Operational Definitions
In this study, polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups. We also classify the types of polarization as follows:
  • Political/ideological polarization: This type of extremism focuses on division, intolerance, and conflict between political parties and followers.
  • Racial or ethnic polarization: This type of polarization focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
  • Religious polarization: This type of polarization focuses on religious identity and incites division, intolerance, and conflict between religious followers
  • Gender polarization: This type of polarization refers to the exclusion, discrimination, and marginalization of individuals based on their gender.
  • Sexual orientation polarization: This refers to the increasing division and distinction between different sexual orientations within society, often leading to heightened tensions, misunderstandings, conflicts, or marginalization among various groups.
  • Other: Polarized texts targeting other groups/identities such as economy class, technology, media, polarization, etc are classified as others.
Furthermore, we define a polarized text manifestations or tactics if it exhibits one or more of the following characteristics:
  • Stereotype: A generalized belief that attributes specific characteristics to all members of a group, often neglecting individual differences, thereby reducing complex personalities to simplistic and uniform representations.
  • Vilification: The act of defaming or demonizing a particular group, person, or entity by inciting fear, often through exaggeration, misrepresentation, or biased framing that portrays the subject negatively and harmfully.
  • Dehumanization: The process of depriving a group or individual of their human qualities or personality by comparing them to animals, machines, or objects, or otherwise denying their humanity, dignity, or individuality.
  • Extreme Language and Absolutism: The use of language that is extreme or makes definitive, all-encompassing statements, often involving words like "always," "never," "worst," or "best," and presenting issues in a dichotomous manner such as "us versus them" or "right versus wrong.
  • Lack of Empathy or Understanding: The absence of compassion or recognition for other viewpoints or experiences in the text.
  • Invalidation: The act of denying or dismissing the identity and existence of individuals or groups, thereby rejecting their sense of self and their presence. See Appendix A for further details and examples for each manifestation of polarization.
3.2 Data Collection
We collected data from various online platforms, including X (formerly Twitter), Facebook, Reddit, Bluesky, Threads, and news/commentary forums. We use a dynamic keyword-driven strategy tailored for each language. Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events. Table˜1 illustrates the compositions, size, and event coverage of the dataset.
The Amharic dataset focuses on the Northern Ethiopia or Tigray War in Ethiopia; Arabic texts cover a broad array of social and economic topics; English data centers on US elections and international conflicts; the German corpus features election discourse and migration debates; the Hausa subset captures religious and ideological discussions in West and Central Africa; Spanish content spans abortion, migration, and indigenous/gender rights; and Urdu texts reflect political and sectarian divides.
Language Source(s) Event Train Dev Test Total Instances
Amharic Facebook, X Tigray War in Northern Ethiopia 3500 500 1000 5000
Arabic Facebook, X, Threads, News Social and economic issues 1482 212 424 2118
English X, BlueSky, News US elections and international conflicts 2117 303 605 3025
German X, BlueSky, Reddit Election discourse and migration debates 2426 347 694 3467
Hausa Facebook, X Religious and ideological discussions in Africa 3893 557 1113 5563
Spanish X, BlueSky Abortion, migration, and indigenous/gender rights 1400 201 401 2002
Urdu X Political and sectarian conflicts 1960 280 560 2800
Table 1: Summary of dataset composition: sources, sizes, and event coverage by task and language.
3.3 Annotation Process
Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators. For crowdsourced annotators, we first filtered candidates based on their previous annotation experience and annotation agreement. Specifically, we selected those who had completed prior tasks with an agreement score of at least 0.8. We also conducted a pilot task consisting of 50 items to further filter annotators based on their agreement scores and average labeling time. For community annotators, we provided a training session followed by a pilot task to assess their performance before assigning the next batch of annotation tasks. During the annotation process, we continuously monitored annotator performance in a batch-wise manner to ensure that the overall process remained on track.
Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language. The English annotation guidelines are provided in Appendix A. Annotators were instructed to:
  • Identify whether a text is polarized;
  • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other);
  • If the text is classified as polarized, tag its manifestations (stereotyping/deindividuation, vilification, dehumanization, extreme language, lack of empathy, invalidation).
Multiple labels were allowed due to the conceptual and contextual overlap often observed in polarized content.
Annotator Recruitment and Annotation Process: We used both expert and crowd-sourced annotation strategies. Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific. To evaluate inter-annotator agreement, we report both Cohen’s Kappa and Fleiss’ Kappa. Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
  • Amharic: We used trained annotators to annotated 5,000 samples in five batches using the POTATO annotation tool. Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control. The annotation quality achieved a Fleiss’ Kappa of 0.49, with full agreement in 69.62% of cases and a peak pairwise Cohen’s Kappa of 0.65.
  • Arabic: Three topic-aware annotators labeled 2,118 samples using POTATO. Final polarization distribution was 16% positive, with Cohen’s Kappa at 0.296. Annotations were revised to ensure thematic consistency.
  • German: Annotation was conducted through Prolific. Annotators were screened via surveys and underwent pilot testing. Final annotation used a locally hosted POTATO interface adapted for multilabel inputs. Inter-annotator agreement was moderate, with notable consistency on polarized cases.
  • Hausa: We used trained annotators and Label Studio with iterative feedback and internal consistency checks. Data focused on religion and ideology on social media.
  • Spanish and English: Amazon Mechanical Turk (MTurk) was used with strict quality filters, such as task completion history and approval ratings. Annotators received feedback and were contacted throughout for quality assurance. English annotation reached a Kappa of 0.52 (up from 0.31 after worker filtering); Spanish showed lower agreement with an average Kappa of 0.24.
  • Urdu: Two approaches were used: (1) Manual Annotation: 1,792 samples labeled in batches with average Cohen’s Kappa of 0.55 (batch 1: 0.79; batch 2: 0.56; batch 3: 0.39). Disagreements were resolved by a fifth annotator. (2) Prolific: 1,460 samples were annotated in three batches. Initial agreement was low (Fleiss’ Kappa = 0.05). Additional filters (education, region) raised agreement in later batches to 0.28 and 0.30. The final dataset contains 2,800 high-quality Urdu samples.
Figure 3: Percentage of polarization per language
3.4 Dataset Characteristics
Language Total Polarization Polarization Types Polarization Manifestations
gender/sexual political religious racial/ethnic other vilification extreme_language stereotype invalidation lack_of_empathy dehumanization
Amharic 5000 3753 29 3342 99 1297 189 2398 1527 2729 799 880 657
Arabic 2118 338 102 178 86 171 220 413 341 283 240 175 102
English 3025 1021 18 892 44 98 5 342 179 138 85 45 39
German 3467 1596 241 1981 448 782 653 1371 1001 1420 1258 1124 568
Hausa 5563 625 46 285 149 183 22 74 179 253 13 53 202
Spanish 2002 1057 254 742 571 521 405 798 622 843 205 599 307
Urdu 2800 1860 236 1533 557 423 126 1548 1409 979 670 632 501
Table 2: Number of samples labeled positive for each annotation task across languages. Labels are grouped by task: Polarization (Task 1), Polarization Types (Task 2), and Polarization Manifestations (Task 3).
We now analyze key trends and insights across the dataset based on annotated labels. Table 2 provides a detailed quantitative breakdown of positive labels for each annotation task.
Polarization Prevalence: As shown in Figure 3 and further detailed in Table 2, Amharic (75%) and Urdu (66%) have the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate. Spanish (53%), German (46%), and English (34%) fall in the mid-range. Arabic (12%) and Hausa (11%) reflect broader topical focus and possibly platform moderation effects. Hausa shows the lowest polarization rate despite covering sensitive religious and ideological topics, possibly due to moderation or cultural norms in expression. Datasets with acute sociopolitical events (e.g., Amharic, Urdu) show higher polarization. Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
Types of Polarization: As illustrated in Figure LABEL:fig:types and Table 2, political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%). Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German. Gender/sexual identity polarization is generally low but visible in Spanish (13%) and Urdu (8%). Thematic emphasis (e.g., migration, elections) shapes type distribution. Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
Polarization Manifestations: Figure LABEL:fig:manifestation and Table 2 show the types of manifestations. Stereotyping/deindividuation and vilification are most frequent in Urdu, Amharic, German, and Spanish. Urdu’s vilification rate is highest (55%). Extreme language is strong in Urdu (50%), Spanish (31%), and Amharic (31%). Invalidation and dehumanization peak in German and Urdu. Conflict-heavy datasets (Urdu, Amharic) correlate with higher rates of hostile manifestations. Platform moderation and topic scope likely suppress such features in Arabic and Hausa.
Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization. The POLAR dataset offers a robust benchmark for examining these effects at scale.
4 Experimentation and Results
4.1 Experimental Setup
To evaluate POLAR , we conducted baseline experiments on three polarization detection tasks:
  • Task 1: Binary classification - determining whether a text is polarized.
  • Task 2: Multi-label classification - identifying polarization types (e.g., political, religious, ethnic).
  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., stereotyping, vilification, dehumanization).
For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table˜1. We pursued two main experimental paradigms:
  1. 1.
    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3) for monolingual and crosslingual experiments.
  2. 2.
[…]
4.2 Results and Analysis
Monolingual Cosslingual
Task Lang. InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa
1. Polarization Amharic 77.96 78.92 81.93 80.64 53.71 74.98 23.26 49.44 5.63 0.00 0.00 83.21
Arabic 58.39 61.94 67.19 70.42 52.48 54.68 29.68 35.87 27.76 2.78 37.79 25.63
English 72.09 69.84 75.43 76.08 72.77 74.94 1.90 51.68 53.41 46.89 33.44 53.45
German 29.75 63.46 67.50 67.78 58.45 64.38 0.62 51.97 64.59 41.87 46.05 25.94
Hausa 59.57 66.41 67.43 66.41 59.32 65.62 8.21 22.92 6.47 1.50 6.25 25.13
Spanish 38.69 64.04 70.98 57.00 59.31 54.16 45.70 68.66 67.26 21.09 68.34 62.17
Urdu 1.07 66.35 79.95 43.91 65.63 60.68 1.60 30.36 0.00 0.00 3.68 68.67
2. Types Amharic 24.22 38.13 43.65 25.71 17.93 25.00 11.56 20.65 2.49 0.00 0.00 26.76
Arabic 22.97 40.23 42.12 37.58 27.53 35.22 11.49 23.73 8.78 2.11 12.47 7.93
English 16.04 23.25 31.38 24.10 21.07 17.70 10.84 19.69 16.42 12.22 11.70 3.93
German 13.52 58.58 61.19 58.56 54.13 40.64 12.03 35.86 29.59 9.72 23.98 11.88
Hausa 18.56 19.14 17.38 18.09 19.61 18.87 3.20 9.97 3.91 1.39 5.78 5.90
Spanish 43.26 66.07 67.76 58.07 57.94 43.40 7.89 47.06 15.38 0.43 32.02 14.60
Urdu 27.14 51.94 51.60 45.38 38.02 33.13 3.77 13.62 6.33 0.00 3.94 20.30
3. Manifestations Amharic 43.52 47.56 47.63 43.17 33.18 43.29 15.57 27.07 8.64 0.00 0.00 43.58
Arabic 40.05 51.55 52.52 55.61 42.18 47.73 16.56 30.68 22.14 0.00 19.57 17.24
English 14.40 15.01 19.39 18.61 18.60 15.15 7.16 10.16 10.05 5.62 10.69 8.85
German 38.49 49.88 52.74 51.70 46.85 51.91 2.38 36.12 27.05 0.00 23.38 12.93
Hausa 19.23 20.04 19.18 18.89 18.74 18.93 5.74 5.86 3.77 3.12 6.19 6.43
Spanish 38.94 50.00 51.04 45.02 45.17 35.09 2.34 40.63 11.83 0.40 35.99 23.56
Urdu 34.26 52.20 53.64 41.32 45.90 47.01 2.10 19.16 11.54 0.00 1.88 48.58
Table 3: Average F1-Macro for all three tasks. In the monolingual settings, we train and evaluate the model for each language separately. In the crosslingual setting, we train on all languages within a language family (AfroAsian: Amharic, Arabic, Hausa, and Urdu), Germanic: English, German, and Spanish) except the target language, and evaluate on the test set of the target language. The best performance scores are highlighted in blue and orange, respectively.
4.2.1 Monolingual Setup Results
To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3). Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
For Task 1, binary polarization classification, RemBERT consistently achieved the highest macro F1-scores for most languages, while XLM-R performed best in Arabic, English, and German. In the multi-label polarization types classification (Task 2), RemBERT again provided the top scores for Spanish, except for Hausa and Urdu, where mBERT and LaBSE performed best, respectively. For Task 3, multi-label polarization manifestations classification, the results were generally lower across all models, with RemBERT delivering the highest scores for most languages, except for Arabic and Hausa.
4.2.2 Cross-Lingual MLM Results
For cross-lingual transfer experiments, each MLM was trained on all other languages within the same language family (Afro-Asian or Germanic), excluding the target language, and evaluated on the target language’s test set for the same three tasks.
In the cross-lingual setup for polarization classification (Task 1), mDeBERTa achieved the highest F1-macro for Amharic, English, Hausa, and Urdu. For Arabic, German, and Spanish, mBERT, RemBERT, and LaBSE performed slightly better than the other models, respectively.
For Task 2, polarization types classification, performance was generally lower as expected in the cross-lingual scenario. LaBSE yielded the best results for four languages, while mDeBERTa performed better for Amharic and Urdu.
Similarly, for Task 3, polarization manifestation classification, the results were quite low overall, but LaBSE and mDeBERTa showed relatively better performance compared to the other models.
This suggests that general polarization detection transfers more easily across languages than the more nuanced identification of polarization types and manifestations, which are culturally and linguistically specific. Effective cross-lingual transfer likely requires aligned annotation schemes and fine-tuning strategies sensitive to these differences.
4.2.3 Zero and Few-shot LLM Performance
We evaluated seven state-of-the-art LLMs (Qwen3-8B, LLaMA-3.1-8B, Mixtral-8x7B, GPT-4o, GPT-4o-mini, GPT-4.1-mini, and Llama-4-Scout-17B-16E-Instruct) across languages. The prompts used for evaluation are provided in the Appendix B. We explicitly provided definitions of tasks and leveraged the models’ chain-of-thoughts capabilities by encouraging reasoning before labeling. In the few-shot setting, we provided 4 example instances for each task. To assess the impact of prompt language, we also compared model performance when prompted in English versus the respective target language.
As Table  4 illustrates, the closed-source models GPT-4o, GPT-4o mini, and GPT-4.1 mini, consistently outperform the other four LLMs, demonstrating superior performance in polarization recognition. However, their performance on Hausa remains notably low, highlighting both the challenge of understanding low-resource languages and the value of our multilingual dataset in exposing the limitations.
In contrast, smaller open-source models, such as Qwen3-8B, LLaMA-3.1-8B, and Mixtral-8x7B, achieve lower F1-Macro scores and perform inconsistently across languages. Their F1-Macro scores improve by 5 10% when provided with few-shot examples. However, the larger LLaMA model, LLaMA-4-scout-17B, demonstrates performance comparable to GPT-4o in polarization classification.
An interesting phenomenon is observed: Smaller models show improved classification performance in the few-shot setting, while larger models such as the GPT series and LLaMA-4-17B do not consistently gain additional improvement with few-shot examples. This suggests that few-shot learning offers advantages to smaller models than to larger ones.
Task Lang GPT 4o GPT 4o-mini GPT 4.1-mini Mistral-7B Qwen3-8B Llama-3.1-8B Llama4-17B
Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot
1. Polarization Amharic 71.30 68.74 55.40 58.04 63.52 63.20 44.84 48.24 55.60 65.51 50.98 58.84 69.96 67.91
Arabic 71.51 79.65 58.21 68.41 76.20 79.58 33.65 53.06 65.95 70.76 61.51 67.21 73.58 77.45
English 75.46 79.56 64.78 80.30 77.43 77.48 53.78 70.00 69.09 66.86 66.55 69.90 76.90 75.48
German 72.04 71.12 63.03 68.49 73.77 69.60 56.17 61.77 65.42 68.76 59.90 65.40 70.89 70.35
Hausa 34.73 44.87 28.21 34.07 39.45 48.06 19.82 28.15 45.91 49.18 38.73 40.80 47.85 54.53
Spanish 72.69 65.76 62.51 67.96 71.28 69.22 51.24 61.00 65.42 68.76 60.20 62.53 66.11 66.27
Urdu 72.25 78.83 59.19 69.92 73.15 72.63 48.27 54.79 66.99 70.89 55.61 59.77 73.93 68.50
2. Types Amharic 39.67 36.59 33.8 34.89 32.49 34.44 44.21 63.24 63.07 75.90 78.26
Arabic 35.00 39.30 37.18 36.82 33.07 36.77 49.39 53.91 53.22 52.86 55.03
English 34.96 35.67 33.79 38.73 35.22 38.22 48.28 68.50 67.37 62.71 67.09
German 58.86 58.58 54.36 57.94 52.70 55.15 44.10 71.82 70.62 73.89 76.20
Hausa 18.96 19.11 16.90 18.15 17.00 18.86 43.34 67.24 70.47 71.75 74.66
Spanish 48.98 52.93 52.82 55.09 48.89 49.76 43.15 65.96 67.87 69.82 71.43
Urdu 55.6 59.64 53.97 54.78 51.82 55.09 48.40 63.36 64.83 60.01 61.52
3. Manifestations Amharic 44.09 45.59 40.01 44.78 40.53 42.49 0 0 0 0 0
Arabic 51.90 56.90 46.96 51.01 52.57 60.76 0 0 0 0 0
English 22.95 25.87 20.43 25.78 28.41 26.67 0 0 0 0 0
German 42.84 45.16 43.33 53.78 37.62 41.23 0 0 0 0 0
Hausa 8.58 8.14 8.42 8.37 8.55 9.33 0 0 0 0 0
Spanish 41.00 40.99 39.85 46.33 38.82 40.66 0 0 0 0 0
Urdu 52.89 54.43 48.25 58.17 48.34 51.89 0 0 0 0 0
Table 4: F1-Macro resulting from the zero- and few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
5 Discussion
Our multilingual and cross-cultural investigation into online polarization reveals that polarization is a deeply contextual and event-driven phenomenon. Its prevalence and intensity are closely tied to the sociopolitical climate, public discourse, and digital environment of each region. In societies experiencing major crises, such as the war in Ethiopia, religious tensions in Pakistan, and international conflicts like the wars in Ukraine and Gaza that influence the US election, we find polarization to be especially amplified on online platforms. Conversely, broader topical selection, moderation, or more diffuse societal tensions are associated with noticeably lower polarization.
The types of polarization evident in online discourse are shaped by the most salient local issues: political (e.g., in Ethiopia, the US, and Germany), religious (e.g., in Pakistan), ethnic (e.g., in Ethiopia), or gender-based (e.g., in Spain) divides tend to dominate where these subjects are at the forefront of public debate or conflict. This underscores the importance of accounting for each region’s unique sociopolitical history and cultural landscape when analyzing or intervening in digital polarization.
Polarization also manifests in varied rhetorical forms, including stereotyping/deindividuation, vilification, and exclusionary or hostile language. These patterns are particularly pronounced amid contentious issues and persistently divided communities, such as civil war, election, migration issues, religious divide, and so on.
From a computational perspective, our results demonstrate both the feasibility and the challenges of detecting polarization in multilingual settings. Binary polarization classification is relatively tractable for well-resourced languages and clear-cut contexts, especially when leveraging state-of-the-art multilingual models (Sec. 4). However, classifying the specific types and manifestations of polarization introduces significantly greater ambiguity and requires a much deeper, context-sensitive understanding. Performance on these more nuanced tasks remains limited. This highlights important directions for future work, including the integration of cultural signals and context or event understanding, the development of more robust multilingual embeddings, and the pursuit of consistent annotation (with a great deal of annotator training and follow up, specially for the crowdsourcing annotations). The wide variability across languages, contexts, and events further affirms the necessity of designing language, culture, and event specific approaches, particularly for underrepresented regions in global NLP research.
Taken together, these findings highlight that effective detection and mitigation of online polarization must move beyond generic or monocultural solutions. There is a pressing need for context-aware, adaptable methodologies that acknowledge both the universal and the local characteristics of polarized discourse, ensuring relevance and effectiveness across the world’s diverse digital landscapes.
6 Conclusion
In this study, we introduced POLAR, the first multilingual, multicultural, and multievent dataset for benchmarking online polarization. Our findings showed that polarization was deeply context-dependent, manifesting in diverse types and rhetorical forms shaped by local sociopolitical dynamics. Theoretically, we provide new evidence that polarization is a multifaceted, cross-lingual phenomenon strongly influenced by language, culture, and current events, rather than being uniform across settings. On a practical level, our benchmark revealed that while binary detection of polarization was feasible with current models, accurately identifying specific types and manifestations remained a persistent challenge, especially for multilingual and low-resource contexts.
This work established a foundation for further computational studies of digital polarization and for developing more culturally robust moderation and intervention tools. Looking ahead, future research should integrate richer contextual and cultural signals into model architectures, refine annotation guidelines for cross-regional consistency, and expand evaluations to include additional languages, social domains, and further events. By releasing our data and benchmarks, we aimed to catalyze further innovation toward nuanced detection and effective mitigation of online polarization worldwide.
Limitations
While POLAR represents an important step toward multilingual, multicultural, and multievent polarization analysis, several limitations remain. First, annotator understanding - particularly in crowdsourced setups - was sometimes limited, potentially impacting label quality. We mitigated this through strict quality assurance methods, including control questions, pre-study surveys, and ongoing annotator assessment, but some variability in interpretation may persist.
Second, in-house annotation, while yielding higher consistency, sometimes introduced psychological challenges for annotators given the sensitive or hostile nature of polarized content. To address this, we provided detailed instructions and support resources to reduce stress and clarify expectations, but some emotional burden may have remained.
Third, our choice of models is not exhaustive. Although we included several leading multilingual models and both open and closed LLMs. Adding more language-specific models in the future could improve results, especially for monolingual scenarios.
Finally, for some of the languages in our benchmark, the available data size is still limited, which may constrain the generalizability of model training and evaluation for those cases. Future work should expand dataset size and diversity, and explore language- or region-specific model development to better support underrepresented contexts.
Ethics Statement
This research uses only publicly available, anonymized data and addresses sensitive topics around polarization in diverse cultures. All annotation was conducted by native speakers using culturally appropriate guidelines; annotators were informed of the project’s social good aims, possible distress, and could opt out anytime. Annotators received prompt and fair compensation above local wage standards or per Prolific’s requirements. Despite rigorous protocols, labeling polarization remains subjective; we encourage responsible, ethically grounded use of this resource and discourage misuse.
As Table 5 demonstrated, the inter-annotation agreement scores range from 0.2 to 0.6, corresponding to fair to moderate levels of agreement according to established benchmarks (landis1977kappa). This level of agreement further highlights the inherent difficulty of polarization recognition and underscores the necessity and value of our proposed work.
Language Annotation Platform Kappa Score
Amharic Potato 0.65
Arabic Potato 0.296
English MTurk 0.52
German Prolific 0.29
Hausa Potato 0.48
Spanish MTurk 0.24
Urdu Potato 0.55
Table 5: Dataset annotation platforms and inter-annotator agreement scores
Appendix A Annotation Guidelines
Annotation Guidelines
To minimize potential misunderstandings, we offer annotation guidelines customized for each event, accompanied by event-specific examples. These guidelines are in the languages of the target datsets. The English version is provided below.
This guideline aims to assess whether social media messages reflect attitude polarization and to categorize the various types and manifestations of this polarization. The dataset consists of content sourced from platforms such as Facebook and Twitter.
Warning to annotators: You may encounter polarized and hateful content during this task. Please take breaks as needed and prioritize your mental well-being.
If you choose to participate, please read the following sections carefully. To exit the study at any time, click “Cancel” or “Quit” and provide feedback if possible.
Task 1: Polarization Classification (Yes/No)
Polarization refers to increasingly extreme, divided beliefs or behaviors between opposing groups. Attitude polarization includes:
  • Negative attitudes toward out-groups
  • Blind support for in-groups
  • Stereotyping, vilification, dehumanization, or intolerance
Texts should be labeled:
  • Yes – if the message clearly reflects attitude polarization
  • No – if it does not show any polarization indicators
Note: Always consider the overall context and meaning, not just individual words.
Task 2: Type of Polarization (if “Yes” in Task 1)
If the message is polarized, select all applicable types:
  1. 2.1
    Political/Ideological Polarization – Conflict between political parties or ideologies
  2. 2.2
    Racial/Ethnic Polarization – Division based on race or ethnicity
  3. 2.3
    Religious Polarization – Conflict based on religious beliefs
  4. 2.4
    Gender Polarization – Discrimination based on gender
  5. 2.5
    Sexual Orientation Polarization – Tension or marginalization based on sexual orientation
  6. 2.6
    Other – e.g., based on economic class, technology, media, etc.
Note: Multiple types may be selected if applicable.
Task 3: Manifestations of Polarization (if “Yes” in Task 1)
Classify the tactics used to express polarization in the message. A text is considered polarizing if it includes one or more of the following:
  1. 3.1
    Stereotyping – Generalizing traits to all group members Example: “Women are weak.”
  2. 3.2
    Vilification – Defaming or demonizing a group Example: “Migrants are traitors.”
  3. 3.3
    Dehumanization – Stripping away human qualities Example: “They are cockroaches.”
  4. 3.4
    Deindividuation – Ignoring individual identities Example: “All migrants are thieves.”
  5. 3.5
    Extreme Language or Absolutism – Using words like “always,” “never,” or “worst” Example: “We can never trust them.”
  6. 3.6
    Lack of Empathy or Understanding – Dismissing others’ perspectives Example: “Wearing the hijab is extremist.”
  7. 3.7
    Invalidation – Denying a group’s identity or existence Example: “There is no nation called Palestine.”
Thank you for your valuable contribution to this research.
Appendix B Prompt for Text Classification
Task 1: Polarization Detection – Zero Shot
Given the following texts in [X] language, classify them as polarized or non-polarized based on the following definition:
Definition of Polarization: Polarization refers to the phenomenon where opinions, beliefs, or behaviors become more extreme or divided, resulting in increased separation and conflict between different groups. These groups may be based on politics or ideology, ethnicity or race, religion, gender, or sexual orientation. Polarization involves negative attitudes such as stereotyping, denigration, dehumanization, de-individualization, or intolerance toward individuals or groups outside one’s own, while exhibiting blind support and solidarity toward those within the group.
Reasoning: Provide the reason why the text demonstrates polarization.
Label: Return True if the text is polarized, otherwise False. Do not return any other value.
Task 1: Polarization Detection – Few Shot
Same instructions as above. Examples provided:
  • Text: Black immigrants should leave the USA.
    Reasoning: This statement stereotypes a group based on race and immigration status, implying a negative attitude and exclusion.
    Label: True
  • Text: Africans are lazy because they are backward.
    Reasoning: Negative trait assignment based on ethnicity, showing polarization.
    Label: True
  • Text: Muslims are terrorists.
    Reasoning: Promotes harmful stereotypes and dehumanization.
    Label: True
  • Text: African women are beautiful.
    Reasoning: Positive generalization, not polarizing.
    Label: False
Task 2: Polarized Text Typing – Zero Shot
Classify each polarized text into one or more of the following types:
  • Gender/Sexual: Discrimination based on gender or sexual orientation
  • Political: Division or conflict between political ideologies
  • Religious: Conflict between religious groups
  • Racial/Ethnic: Conflict or division based on race or ethnicity
  • Other: Related to economy, technology, media, etc.
Output: A list of five binary values in the order: [Gender/Sexual, Political, Religious, Racial/Ethnic, Other]
Few Shot Examples
  • Text: Black immigrants should leave the USA.
    Output: [0, 0, 0, 1, 0]
  • Text: Muslims are terrorists.
    Output: [0, 0, 1, 0, 0]
  • Text: Women should be confined to traditional roles.
    Output: [1, 0, 0, 0, 0]
  • Text: The economy is failing because of government policies.
    Output: [0, 1, 0, 0, 0]
Task 3: Polarization Manifestations – Zero Shot
Analyze and identify the following manifestations in each text. Output a list of six binary values in this order:
[Vilification, Extreme Language, Stereotype, Invalidation, Lack of Empathy, Dehumanization]
Definitions
  • Stereotype/Deindividuation: Generalizes traits, ignores individuality.
  • Vilification: Defames a group or person, inciting fear.
  • Dehumanization: Strips humanity, compares people to animals or objects.
  • Extreme Language: Uses absolutes like always, never, us vs. them.
  • Lack of Empathy: Dismisses others’ experiences or views.
  • Invalidation: Denies the identity or existence of a group.
Few Shot Examples
  • Text: Black immigrants should leave the USA.
    Output: [1, 1, 1, 1, 1, 0]
  • Text: Africans are lazy because they are backward.
    Output: [1, 1, 1, 1, 1, 0]
  • Text: Muslims are terrorists.
    Output: [1, 1, 1, 1, 1, 1]
  • Text: African women are beautiful.
    Output: [0, 0, 0, 0, 0, 0]
Appendix C Dataset Samples
Table 6: Examples of polarized statements with annotated types and manifestations. Types: P: Political, Re: Religious, Ra:Racial; Manifestations: V: Vilification, EL: Extreme Language, S: Stereotype, I: Invalidation, LE: Lack of Empathy, D: Dehumanization.
Text T2 T3
"Denazification" as an excuse to explain Russia’s deployment of troops to Ukraine, alleging that Hitler was of Jewish ancestry P V, EL
"Heck, I voted for Biden three times." Signed, Liberal Asshats in Swing States P EL, S, D
Satellite images confirm civilian massacres in towns during Russian military occupation P V
"Sufficient evidence of voter fraud?" Where? marjorieisaninsurrectionist P EL, I
"They" are now the deep state. P S, I, D
15000 dead children claimed by Hamas who uses those children as human shields. Yes yes we covered that already. Now you’re just talking in circles P,Ra V, EL, I, LE, D
5 Black Officers Awarded \$16M After White Colleague’s Racial Comments R V
A Federal Investigation needs to be launched against TWITTER Fraud and Election Interference P EL, I
A Jewish state was created because no one else wanted them. This is why. Zionism and Naziism are the same disease. P+Re+Ra V, EL, S, I, LE, D
A lot of lying by the radical left. P V, S, I, D
A rational and just society probably wouldn’t allow school shootings to be part of their National identity P EL, S, I, D
A relationship between the Apartheid state and a tiny tyrant state. Free Palestine P V, I, LE, D
A small price to pay for Ukrainian sovereignty and our green future. Stop whining. P I, LE
A traitor today, a traitor tomorrow, a traitor always! P V
A valid protest of a rigged election, that ended up having some sort of kerfuffle. P V, I
A very moral army. IDF Gaza Israel P V, LE
Text Lang Label
Gibt sogar recht viele Ex-Muslime. Alles sehr vernünftige Leute. DE TRUE
There are actually quite a few ex-Muslims. All very reasonable people. EN
Kriegstreiber! Auch ihr, die Propagandapresse!!! DE TRUE
Warmongers! You too, the propaganda press!!! EN
Steuerzahler und Juden DE TRUE
Taxpayers and Jews EN
Ja die Türken leiden sehr stark unter ihrem links grün versifften Erdogan. Muss man wissen. :D DE TRUE
Yes, the Turks are suffering greatly under their left-green filthy Erdogan. Good to know. :D EN
Die Klima Armee Fraktion wird kommen. DE TRUE
The Climate Army faction will come. EN
Denk ich nicht, nein. Migration ist deren Kernthema. DE TRUE
I don’t think so, no. Migration is their core issue. EN
"Es gibt bereits vereinzelte Einsätze von Bodentruppen in Gaza. @URL" DE FALSE
"There are already isolated deployments of ground troops in Gaza. @URL" EN
Lass dich nicht aufhalten. Es gibt eine ukrainische Fremdenlegion. Da kannst du dich melden. DE FALSE
Don’t let yourself be stopped. There is a Ukrainian foreign legion. You can sign up there. EN
Klingt für mich nach Long COVID DE FALSE
Sounds like long COVID to me. EN
Wenn ich schon wieder dieses Woke lese. Der AfD ist sogar unser Grundgesetz zu Woke. Hört auf das Wort so inflationär zu gebrauchen. DE FALSE
When I read this "woke" again. Even our constitution is too woke for the AfD. Stop using the word so inflationarily. EN
CO2 ist gut für Pflanzen DE FALSE
CO2 is good for plants. EN
Table 7: German samples with labels and English translations
Text Lang Label
y vos un random judio que comenta pelotudeces para poder figurar en algun lado SPA TRUE
And you, a random Jew who comments nonsense just to get noticed somewhere. EN
quiero comer indio viejo SPA TRUE
I want to eat old Indian. EN
es que a dia de hoy parece que la diferencia es nazis con aborto y nazis sin aborto SPA TRUE
Nowadays it seems that the difference is Nazis with abortion and Nazis without abortion. EN
orgulloso de sacarle la conchetumare a un pastor evangelico por meterse en weas que no debia. SPA TRUE
Proud to beat the hell out of an evangelical pastor for meddling in things he shouldn’t. EN
el tribunal superior del partido corrupto SPA TRUE
The supreme court of the corrupt party. EN
claudia es judia, que no manche SPA FALSE
Claudia is Jewish, don’t stain. EN
la escuela feminista de pintura SPA FALSE
The feminist school of painting. EN
la cancion que me representa ahora que estoy deportado de los estados unidos. SPA FALSE
The song that represents me now that I am deported from the United States. EN
aunque, siendo justos, esa foto es de una deportacion de 2018, cuando trump estaba gobernando. SPA FALSE
Although, to be fair, that photo is from a 2018 deportation, when Trump was in power. EN
la propuesta de k!nadim me parece muy interesante de cara a eurovision, es facilona y divertida #benidormfestsemi1 SPA FALSE
The proposal by k!nadim seems very interesting to me for Eurovision; it’s easy and fun. #benidormfestsemi1 EN
Table 8: Spanish sentences with labels and English translations.